OpenStudy (anonymous):

ABCD is a parallelogram. Segment AC = 4x + 10. Find the value of x and y.

OpenStudy (anonymous):

pic

OpenStudy (anonymous):

i dont know if i am doing this correctly but with a parallelogram ae=ec so 2 x ae or 2x ec would equal ac right?

OpenStudy (anonymous):

sounds good to me thank you

OpenStudy (anonymous):

im not done yet

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

but is that all you want

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

nope..lol

OpenStudy (anonymous):

okies

OpenStudy (anonymous):

we know this (3x+y)+(2x+y) = 4x+10 haha dumb me why didnt i think of that before

OpenStudy (anonymous):

then solve it like this :P

OpenStudy (anonymous):

awesome thank you

OpenStudy (anonymous):

its not done but thats the beginning do you know how to continue

OpenStudy (anonymous):

so x is o and y is 10 right

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

x=0 is correct, however for y we have, 3x+2x+y+y=4x+10 we know x=0 so 2y=10 ---> y=5

OpenStudy (anonymous):

(0,10)